OpenStudy (anonymous):

the area of a regular octogon is 35cm squared. what is the area of a regular octogon with sides four times as long medals rewarded for each answer

OpenStudy (nurali):

\[(\frac{ 1 }{ 4 })^2=\frac{ 35 }{ x }\]

OpenStudy (nurali):

These two regular octagons are similar. So, the theorem we used on the previous problem holds here. The square of the scale factor of two similar polygons is equal to the ratio of their areas. In words, the proportion is: the square of (one is to four) = 35 to x where x is the area of the octagon that has sides of length four times the sides of the smaller octagon.

OpenStudy (nurali):

\[\frac{ 1 }{ 16 }=\frac{ 35 }{ x }\] Cross Multiply \[x=35*16\]

OpenStudy (nurali):

\[x=?\]

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry i think it's two times

OpenStudy (anonymous):